### v2

# Defense Against Prompt Injection Attack by Leveraging Attack Techniques

Abstract
With the advancement of technology, large language models (LLMs) have achieved remarkable performance across various natural language processing (NLP) tasks, powering LLM-integrated applications like Microsoft Copilot. However, as LLMs continue to evolve, new vulnerabilities, especially prompt injection attacks arise. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in data content, such as retrieved results. Recent attack methods leverage LLMs’ instruction-following abilities and their inabilities to distinguish instructions injected in the data content, and achieve a high attack success rate (ASR). When comparing the attack and defense methods, we interestingly find that they share similar design goals, of inducing the model to ignore unwanted instructions and instead to execute wanted instructions. Therefore, we raise an intuitive question: Could these attack techniques be utilized for defensive purposes? In this paper, we invert the intention of prompt injection methods to develop novel defense methods based on previous training-free attack methods, by repeating the attack process but with the original input instruction rather than the injected instruction. Our comprehensive experiments demonstrate that our defense techniques outperform existing defense approaches, achieving state-of-the-art results.
Defense Against Prompt Injection Attack by Leveraging Attack Techniques
Yulin Chen1thanks: Yulin Chen and Haoran Li contributed equally., Haoran Li211footnotemark: 1, Zihao Zheng3, Yangqiu Song2, Dekai Wu2, Bryan Hooi1 1National University of Singapore 2The Hong Kong University of Science and Technology 3Harbin Institute of Technology, Shenzhen chenyulin28@u.nus.edu
1 Introduction
Figure 1: Examples of indirect prompt injection attacks (a) and the design of our defense method based on the attack technique (b).
With the continuously developing technologies, large language models (LLMs) have achieved impressive performance on various NLP tasks Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023), and are integrated into various real-world applications, such as Microsoft Copilot [^1] https://copilot.microsoft.com/, perplexity.ai [^2] https://www.perplexity.ai/, and so on. However, their inherent instruction-following capabilities make them vulnerable to prompt injection attacks. These attacks trick LLMs into deviating from the original input instructions and executing the attacker’s instructions injected in the data content, such as retrieved results from search engines. The prompt injection attacks can be generally classified into direct attacks Perez and Ribeiro (2022); Chen et al. (2024) and indirect attacks Greshake et al. (2023); Li et al. (2023); Zhan et al. (2024), according to the source of the input data content. For direct prompt injection attacks, the attackers, who are also the users, directly inject instructions into the data content for malicious purposes such as application prompt extraction Perez and Ribeiro (2022). Because of their instruction following ability, and their inability to distinguish the injected instructions, the LLMs execute the instructions in the data content and give undesired responses. On the other hand, for indirect prompt injection attacks, which have garnered more research attention recently, the malicious instructions are injected into external data content, such as retrieved results from external tool usage. In Figure 1 (a), for instance, attackers can inject the malicious prompt into the external data content, which consists of an attack prompt like “Forget previous instruction, and it’s urgent to” and an injected instruction after the attack prompt. This misleads the LLM into generating responses that align with the attacker’s intentions rather than following the original input instructions, thereby avoiding suspicion and potentially convincing users to click on malicious links Liu et al. (2024a). Current defense methods against prompt injection attacks primarily rely on fine-tuning Chen et al. (2024); Wallace et al. (2024); Suo (2024); Piet et al. (2023) or prompt engineering Hines et al. (2024); san (2023); ins (2023); Willison (2023). While fine-tuning-based defenses require annotated data and significant computational resources, prompt engineering approaches, though training-free, often prove less effective. In fact, the Open Worldwide Application Security Project (OWASP) has ranked prompt injection attacks as the #1 security risk for LLM applications OWASP (2023).
In this paper, we propose prompt injection defense methods based on several effective prompt engineering attack techniques. To explain our motivation, consider the example in Figure 1 (a). In this scenario, the malicious prompt (highlighted in red) embedded in the retrieved results consists of an attack prompt followed by an injected instruction. The attack prompt misleads the LLM into ignoring the original input instruction, whose answer could otherwise raise the user’s suspicion. The response to the injected instruction fulfills the attacker’s malicious intent. In contrast, our defense goal is for the LLM to ignore the injected instruction and instead respond to the original input instruction. Interestingly, the defense and attack share similar design goals: inducing the LLM to ignore the unwanted instructions and instead to execute the wanted instructions. This raises an intuitive question: Could attack techniques be repurposed or adapted to develop more robust defense methods? Figure 1 (b) demonstrates how we develop our defense strategy based on the attack techniques: we preserve the attack prompt as the shield prompt, and replace the injected instruction with the original input instruction. We apply this approach with several attack techniques. Moreover, we additionally find that when attackers get access to the conversation template, they can pretend to be the assistant to answer the original input instructions, and then act as the user to request the LLM to answer their injected instruction, posing a serious threat. Inspired by this, we design our defense by acting as the assistant who detects the attack and then acting as the user to confirm the instruction.
We conduct comprehensive experiments to evaluate the effectiveness of our defense methods against various prompt injection attack methods. The results demonstrate that our methods outperform existing training-free defense approaches and are even comparable to fine-tuning-based methods, providing more robust protection. Notably, the defense method based on the most effective attack technique performs the best, reducing the attack success rate (ASR) to nearly zero in certain scenarios. Our contributions are summarized as follows:
  • We present a novel approach to designing defense methods against prompt injection attacks by leveraging effective attack techniques.
  • We develop prompt injection defense methods based on attack strategies, which demonstrate greater effectiveness compared to existing baselines.
  • We significantly reduce the Attack Success Rate (ASR) across various types of attacks, comparing with the previous baselines, with ASR approaching zero in some scenarios.
2 Related Work
2.1 Prompt Injection Attacks
Prompt injection attacks have become a significant challenge for Large Language Models (LLMs), particularly in LLM-integrated applications. These attacks have been widely studied Perez and Ribeiro (2022); Willison (2023); Liu et al. (2023); Li et al. (2023); Liu et al. (2024b); Zhan et al. (2024); Shi et al. (2024); Liu et al. (2024a); Shafran et al. (2024); Huang et al. (2024); Breitenbach et al. (2023). Broadly, prompt injection attack methods can be classified into two categories: prompt-engineering-based attacks Breitenbach et al. (2023); Perez and Ribeiro (2022); Willison (2023); Liu et al. (2024b) and gradient-based attacks Huang et al. (2024); Shafran et al. (2024); Liu et al. (2024a); Shi et al. (2024). In prompt-engineering-based attacks, Perez and Ribeiro (2022) prepend an “ignoring prompt” to the injected instruction, while Willison (2023) propose adding a fake response to convince the LLM that the user’s input has been processed, prompting it to execute the maliciously injected instruction instead. On the other hand, gradient-based attacks, such as those based on the GCG attack method Zou et al. (2023), focus on training a suffix to induce the LLM to produce the desired response.
2.2 Prompt Injection Defenses
Given the severity of prompt injection attacks, several defense methods have been proposed san (2023); Hines et al. (2024); Willison (2023); Chen et al. (2024); Wallace et al. (2024); Yi et al. (2023); Piet et al. (2023); Suo (2024). san (2023) and Yi et al. (2023) suggest appending reminders to reinforce the importance of adhering to the original instructions. Hines et al. (2024) and Willison (2023) propose using special tokens to clearly delineate the data content area. Piet et al. (2023) defend against attacks by training models to perform specific tasks, rendering them incapable of following other potentially malicious instructions. Chen et al. (2024) and Wallace et al. (2024) advocate fine-tuning LLMs with instruction-following datasets, granting privileged status to authorized instructions. Lastly, Suo (2024) introduce a method of signing instructions with special tokens, ensuring that LLMs only follow those that are properly signed.
3 Background
Before introducing our defense methods, we provide an overview of well-known prompt-engineering-based attack techniques, as these form the basis of our defense strategy.
3.1 Naive Attack
The naive attack method involves simply appending the injected instruction to the original data content, as shown in Figure 9. In most cases, the LLMs execute both the original input instruction and the injected instruction, and the response is not misleading or deceptive.
3.2 Escape Characters Attack
[…]
3.3 Ignore Attack
The Ignore attack Perez and Ribeiro (2022) is a commonly used prompt injection attack technique. As illustrated in Figure 12, the attacker crafts an attack prompt that persuades the LLM to disregard the previous instruction and instead to follow the attacker’s injected instruction.
3.4 Fake Completion Attack
As demonstrated in Figure 13, the fake completion attack involves first appending a fake response to the original instruction, misleading the LLM into thinking that the previous instruction has been completed. The attacker then injects their own instruction into the subsequent content.
However, this example represents a relatively weaker attack, as it assumes that the attacker does not have knowledge of the full conversation template. For instance, in the case of Figure 13, the attacker uses ‘###instruction:’ as the instruction identifier, whereas the actual identifier is ‘<Instruction>’. If the attacker has access to the entire conversation template, they can fabricate a more convincing assistant response, as illustrated in Figure 14, making this type of attack much harder to defend against.
4 Methodology
Figure 2: Examples of how we designed our defense methods based on the attack techniques.
4.1 Problem Formulation
Given an input instruction $I$ and clean data content $D$ (which may come from user input or external sources such as search engines), the LLM $M$ generates a benign response $R^{b}$ based on the combination of $I$ and $D$, denoted as $M(I\oplus D)=R^{b}$. In a prompt injection attack, the attacker injects malicious prompt $P$ into the clean data content $D$, causing the LLM $M$ to generate a response $R^{t}$ that reflects the attacker’s intended target, represented as $M(I\oplus D\oplus P)=R^{t}$. To defend against this, we propose a shield prompt $S$ inspired by attack techniques. When $S$ and the original input instruction $I$ are appended to the poisoned data content $D\oplus P$, the LLM still produces the normal response $R^{b}$ without incorporating the attacker’s target, such that $M(I\oplus D\oplus P\oplus S\oplus I)=R^{b}$. Additionally, it is critical that the shield prompt $S$ does not interfere with clean data inference.
4.2 Design Defense from Attack
In Section 3, we introduced prompt-engineering-based attacks. Now, we will explain how we design defense methods inspired by these attack techniques. As described earlier, the attack methods achieve two objectives: 1) tricking LLMs into ignoring the original instruction, and 2) misleading LLMs into executing the injected instruction. These attack methods are highly effective, and we can derive defense strategies from them.
4.2.1 Ignore Defense
The ignore defense is inspired by the ignore attack method. For our defense, the goal is to prevent the model from executing the injected instruction and ensure it follows the original input instruction. The ignore attack strategy serves as a useful guide here. As illustrated in Figure 2 (a), after encountering poisoned data content, we adopt the ignore attack structure by first presenting a shield prompt which is the same as the ignore attack prompt, instructing the LLM to disregard all previous instructions, including both the original and injected ones. We then append the original input instruction to the subsequent content. It’s important to note that the shield prompt can be crafted to be more persuasive than the basic example shown.
4.2.2 Escape Defense
[…]
4.2.3 Fake Completion Defense
Another attack method, the fake completion attack, similarly misleads the model into ignoring the original instruction. In this attack, a fake response (attack prompt) such as ‘### Response: OK’ tricks the LLM into believing that the original instruction has been completed. For our defense, we mimic this approach by repeating the fake response as our shield prompt. As shown in Figure 2 (c), we first fabricate a response to the last instruction. To keep it simple, we use the same response as the attacker’s fake response—‘OK’. This defensive response can evolve alongside improvements in attack techniques. We then append the original input instruction. Upon reading the fabricated response, the LLM will assume the injected instructions have already been executed and will only follow the appended original input instruction. Since the fake completion defense simulates a multi-turn conversation, it could be possible to design the defense within a real multi-turn conversation structure.
4.2.4 Fake Completion Defense with Template
As discussed in Section 3.4, if attackers are aware of the conversation template, they can fabricate assistant responses and create a multi-turn conversation that more convincingly misleads the LLM into believing the original input instruction has already been completed and can be ignored. This motivates us to build our defense within a multi-turn conversation structure. As shown in Figure 2 (d), we first simulate the assistant role and report the presence of a prompt injection attempt (no matter whether true or false). Then the simulated assistant rejects and distrusts all previous instructions, prompting the user to confirm the trusted instruction. Then, we pretend to be the user and confirm the original input instruction.
Defense Methods Llama3-8b-Instruct Qwen2-7b-Instruct Llama3.1-8b-Instruct
Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined
None 46.15 74.51 54.80 64.90 76.92 74.03 85.09 90.86 100.00 100.00 51.92 76.92 62.98 79.80 77.40
Sandwich 21.63 38.46 20.67 18.75 49.51 27.40 47.11 29.80 52.40 67.78 22.59 32.69 22.59 33.17 34.13
Instructional 36.53 35.57 48.07 31.25 29.32 74.03 85.09 83.17 99.03 100.00 39.42 48.55 51.44 62.01 47.11
Reminder 24.51 37.50 36.05 16.82 35.09 78.36 87.01 90.38 99.51 100.00 35.57 56.25 39.42 36.53 42.30
Isolation 37.98 64.90 47.11 62.01 75.48 58.17 73.55 79.80 96.15 98.55 46.63 67.30 59.13 77.88 64.42
Spotlight 27.88 53.36 45.19 75.96 66.34 74.03 78.84 77.40 99.51 99.51 38.94 57.69 41.34 68.75 68.75
Ours-Ignore 11.05 22.11 7.21 7.69 27.40 12.01 11.53 8.65 5.28 16.34 12.50 13.94 5.76 8.17 9.13
Ours-Escape 19.71 38.94 14.90 25.00 34.61 21.63 29.32 16.82 70.19 36.53 12.50 13.94 5.76 8.17 9.13
Ours-Fakecom 16.82 36.53 12.50 0.48 6.25 20.67 13.94 13.46 3.36 6.25 27.40 33.17 22.11 7.21 17.30
Ours-Fakecom-t 11.53 5.28 7.21 0.0 1.44 11.05 7.21 8.17 4.32 2.40 9.13 4.32 3.36 2.40 3.84
Table 1: The results of our defense methods compared to baselines against various attack methods in the direct prompt injection scenario. Bold indicates the best performance. All results are reported in %.
5 Experiments
5.1 Experimental Settings
Dataset.
We evaluate our defense methods against both direct and indirect prompt injection attacks. For direct injection attacks, we follow the method of Chen et al. (2024), applying attacks to 208 samples from AlpacaFarm Dubois et al. (2024) and comparing the effectiveness of our defense methods with baseline approaches. For indirect prompt injection attacks, we use the QA dataset filtered by Li et al. (2023), where malicious instructions are injected into retrieved data content for evaluation, and this dataset contains 2000 samples.
Victim Model.
We select popular and strong open-source LLMs as victim models for our experiments. Specifically, we choose Llama3.1-8b-Instruct Dubey et al. (2024), Qwen2-7b-Instruct Yang et al. (2024), and Llama3-8b-Instruct AI@Meta (2024). Throughout the experiments, unless otherwise specified, “Llama3,” “Llama3.1,” and “Qwen2” refer to Llama3-8b-Instruct, Llama3.1-8b-Instruct, and Qwen2-7b-Instruct, respectively.
Evaluation Metrics.
For the security metric, we follow the evaluation protocol of Chen et al. (2024), using the attack success rate (ASR) to assess the effectiveness of the defense methods. We detecte if the answer to the injected instruction appears in the generated response. For the utility metric, we use accuracy to evaluate the potential negative impact of defense methods on model performance. Specifically, we employ the filtered QA dataset Li et al. (2023), which is not attacked and includes the defense mechanism. We request the LLMs to answer the questions and verify whether the correct (golden) answers appear in the responses.
Defense Methods Llama3-8b-Instruct Qwen2-7b-Instruct Llama3.1-8b-Instruct
Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined
None 10.55 53.35 88.25 75.30 86.00 92.45 95.90 100.00 100.00 100.00 85.90 91.10 81.70 95.25 92.30
Sandwich 0.45 9.35 49.55 7.30 21.25 4.80 6.15 14.00 34.20 34.60 2.50 3.05 22.90 3.35 9.55
Instructional 6.95 35.00 80.10 64.45 62.75 95.55 95.75 99.95 100.00 100.00 60.15 68.35 88.10 84.70 84.85
Reminder 10.55 39.90 67.50 37.85 51.20 97.65 97.95 100.00 100.00 100.00 79.05 77.30 71.75 84.35 80.65
Isolation 2.20 33.75 83.35 67.40 77.75 77.80 88.85 99.35 99.70 100.00 76.75 85.00 89.75 91.70 88.75
Spotlight 8.80 32.85 76.35 74.45 56.60 94.35 96.45 100.00 100.00 100.00 94.35 96.45 100.00 100.00 100.00
Ours-Ignore 0.05 0.35 0.30 0.10 1.35 0.85 0.70 0.80 0.95 4.10 0.25 0.30 0.35 0.45 1.10
Ours-Escape 0.25 1.70 1.05 0.55 1.45 1.45 1.70 0.75 0.68 4.95 1.25 2.70 1.05 0.90 1.65
Ours-Fakecom 0.10 1.80 17.70 0.05 0.10 0.30 0.70 0.55 0.45 0.30 1.75 2.45 8.75 0.80 0.60
Ours-Fakecom-t 0.05 0.05 0.30 0.05 0.05 0.25 0.20 0.15 0.05 0.05 0.05 0.10 0.10 0.05 0.10
Table 2: The results of our defense methods compared to baselines against various attack methods in the indirect prompt injection scenario. Bold indicates the best performance. All results are reported in %.
5.2 Baselines
5.2.1 Attack Methods
As discussed in Section 3, we select the following attack methods for evaluation: Naive attack (abbreviated as “Naive”), Ignore attack (“Ignore”), Escape-Deletion attack (“Escape”), Fake completion attack (“Fakecom”), and Fake completion attack with template (“Fakecom-t”). Additionally, we include a Combined attack Liu et al. (2024b), which combines the Ignore attack, Fake completion attack, and Escape-Separation attack, referred as “Combined.” An example is shown in Figure 15.
5.2.2 Defense Baselines
For a fair comparison, we select existing training-free defense methods as baselines. Specifically, we select Sandwich san (2023), Instructional ins (2023), Reminder Yi et al. (2023), Isolation Willison (2023), Spotlight Hines et al. (2024) for comparison. More details about the baselines can be found in Appendix A.1.
5.3 Resutls and Analysis
5.3.1 Defense against Direct Attack
We perform the direct prompt injection attack following the approach of Chen et al. (2024), using 208 samples from AlpacaFarm. Table 1 presents the effectiveness of our defense methods in the direct prompt injection scenario. The results show that our methods, which are based on attack techniques, outperform the baselines, regardless of the attack method or the victim model. Among the baseline methods, the Sandwich method performs better on average than the others. The key difference between Sandwich and the other baselines lies in the position of the defense prompt: Sandwich places the defense prompt at the end of the data, similar to our methods. This suggests that placing the defense prompt at the end may interfere with the attack and enhance the defense’s effectiveness. When comparing the victim models, we find that Qwen2 is more vulnerable to attacks, possibly indicating that Qwen2 is less capable of distinguishing between different instructions.
5.3.2 Defense against Indirect Attack
In addition to evaluating defense against direct prompt injection attacks, we also assess its effectiveness against indirect attacks. The key difference between direct and indirect prompt injection attacks is that, in the case of indirect attacks, the input data is retrieved from external tools, such as search engines, and users are often unaware of the attack. Typically, the retrieved data is longer than direct user input. To evaluate indirect prompt injection attacks, we use the filtered QA dataset from Li et al. (2023). Table 2 shows the results of our defense methods compared to the baselines in the indirect scenario. Our methods continue to outperform the baselines by a significant margin. When comparing both direct and indirect prompt injection attacks, it appears that indirect attacks are easier to defend against. Furthermore, Qwen2 remains the most susceptible model to attacks compared to the other two models.
5.3.3 Model Utility
A key evaluation metric for defense methods is their potential impact on the model’s utility. To assess the impact of our method, we use the filtered QA dataset from Li et al. (2023). For simplicity, we do not introduce any attacks into the retrieved data content, and we only verify whether the correct (golden) answer appears in the model’s response, with different defense methods. Table 3 presents the utility performance of various defense strategies. Interestingly, most defense strategies do not significantly affect the model’s utility. Moreover, our proposed defense methods can even improve the performance in some scenarios.
Defense Methods Llama3 Llama3.1 Qwen2
None 78.05 77.10 76.60
Sandwich 80.80 79.50 77.35
Instructional 77.30 79.30 75.35
Reminder 77.20 78.05 76.05
Isolation 78.10 78.25 77.10
Spotlight 76.40 77.90 78.25
Ours-Ignore 78.55 79.60 77.60
Ours-Escape 79.40 79.85 80.40
Ours-Fakecom 80.75 80.45 81.30
Ours-Fakecom-t 79.40 80.45 77.40
Table 3: The model utility when applied with different defense methods. All the results are reported in %.
5.3.4 Model Efficiency
Since LLMs are ultimately integrated into various applications, it is crucial to assess the overhead introduced by the defense method, particularly in terms of generating longer, but less informative, responses. To evaluate efficiency, we use the same QA dataset employed for assessing model utility. As demonstrated in Section 5.3.3, the utility results remain almost consistent across different defense strategies. Therefore, any additional overhead indicates an increase in uninformative content within the responses. Table 8 presents the efficiency of various defense methods. The average overhead of our defense method based on "Fakecom-t" and the defense based on escape attack is slightly higher compared to the baseline with no defense, but it remains relatively low. The overhead for the other defense methods is nearly the same.
5.4 Ablation Study
In this section, we address several questions regarding our defense methods and the evaluation metrics. We perform comprehensive experiments to solidify the validity and robustness of our approach.
Can our methods be extended to the closed-source models?
To further validate the effectiveness of our methods, we apply our methods to the closed-source models "GPT-3.5-Turbo" Jiang et al. (2023) and "GPT-4o-Latest" Hurst et al. (2024). Because we cannot change the conversation template, we only compare our methods based on "Ignore attack" and "Fakecom attack" with the baselines against direct prompt injection attack. Table 4 shows the results. From the table we can find out that our methods are also effective on closed-source models, surpassing the previous defense baselines. What’s more, comparing the defense performance of the two models with our defense methods reveals that stronger model is more suitable to our methods, making our methods more applicable.
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
Naive Ignore Escape Fakecom Combined Naive Ignore Escape Fakecom Combined
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Instructional 25.00 34.61 26.92 44.23 71.63 24.51 18.26 28.36 62.01 42.78
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21
Table 4: The results of our defense methods compared with defense baselines applied on closed-source models. Bold indicates the best performance. All results are reported in %.
Can our methods defend against the gradient-based attack?
Beyond prompt-engineering-based attacks, we also examine the effectiveness of our defense methods against the gradient-based attack. We implement direct prompt injection attack with GCG attack method using Llama3, following the implementation by Chen et al. (2024), and construct the suffix with knowledge of the conversation template. Since the GCG suffix is transferable in jailbreak attacks Zou et al. (2023), we also test its transferability in prompt injection attacks using Llama3.1. Table 6 presents the defense results. The first notable observation is that the suffix transferability is limited. Without any defense, the ASR remains relatively low, similar to a naive attack. Comparing with the baselines, our defense methods more effectively mitigate the attack. The method based on “Fakecom-t” is the most effective, reducing the ASR to below 10%.
How effective is the fake completion attack with conversation template?
Although it’s very unlikely for the attacker to be aware of the conversation template—since application providers typically filter out template tokens—we are still interested in assessing the potential harm of such an attack. We utilize the AlpacaFarm for evaluation and table 5 presents the results. The table shows that the fake completion attack with a conversation template can be harmful, and most baseline methods are ineffective. Our methods, which rely on ignoring the attack and using the fake completion strategy, function as intended but result in only a limited decrease in ASR. Our method based on this attack ("Fakecom-t") is effective, and this phenomenon raises our question: Would the effectiveness of the attack methods determine the defense methods designed on them?
Defense Methods Llama3 Llama3.1 Qwen2
None 98.07 99.51 100.00
Sandwich 68.26 53.36 68.26
Instructional 98.07 92.78 100.00
Reminder 97.11 84.13 100.00
Isolation 98.07 100.00 100.00
Spotlight 100.00 100.00 100.00
Ours-Ignore 9.13 23.07 21.63
Ours-Escape 18.26 12.50 50.96
Ours-Fakecom 30.28 50.00 35.09
Ours-Fakecom-t 1.92 16.82 8.17
Table 5: The results demonstrate how harmful the fake completion attack with the conversation template is. All the results are reported in %.
Can a stronger attack method lead to a stronger defense method?
Given the comparative results from previous attack and defense evaluations, we aim to investigate the relationship between the effectiveness of attacks and defenses using the same techniques. For this purpose, we use AlpacaFarm as the evaluation setting. To assess attack strength, we calculate the average ASR across different defense methods, applying the same process to evaluate defense effectiveness. As shown in Figure 3, stronger attacks tend to lead to stronger defenses, with only one exception observed in the Qwen2 and Llama3.1 model. Additionally, the figure reveals that different models exhibit varying levels of vulnerability.
Can our methods compare with the fine-tuning-based methods?
In previous introduction, we argue that fine-tuning-based methods require significant computational resources. But the fine-tuning-based methods are more effective than previous prompt-engineering-based methods. Therefore, we compare our methods with StruQ Chen et al. (2024), which incorporates prompt injection attack methods into the clean data for fine-tuning. We incorporate the "Naive attack" and the "Ignore attack" respectively for evaluation. Table 7 shows the results. It’s obvious that the ability of StruQ to generalize to the unknown attacks is not satisfactory. Because "Ignore attack" is a part of "Combined attack", "StruQ-Ignore" can defend against "Combined attack" successfully. The generalization ability of our methods is much better, effectively defending against different attacks.
Defense Methods Llama3 Llama3.1
None 87.01 47.11
Sandwich 19.23 22.59
Instructional 28.84 37.50
Reminder 24.51 32.21
Isolation 40.38 39.42
Spotlight 19.71 29.80
Ours-Ignore 12.01 11.05
Ours-Escape 19.23 21.15
Ours-Fakecom 13.94 22.59
Ours-Fakecom-t 9.61 6.73
Table 6: The performance of the defense methods against the gradient-based attack. Bold indicates the best performance. All results are reported in %.
5.5 Case Study
Figure 16 provides two examples of responses without defense against the ignore and fake completion attacks, both generated by Llama3. Additionally, we include the response generated by the model with defense method based on "Fakecom-t", as it is the most effective defense approach. From the examples, we observe that the ignore attack does not consistently persuade the LLM to ignore the original input instruction and the model may end up executing both instructions. While in this instance, the fake completion attack successfully misleads the LLM to execute the injected instruction directly, this strategy does not always work—there are cases where the model executes both instructions, explaining the failures of the defense methods based on these attacks. In the case of the defense method based on "Fakecom-t", we can observe that the defense method successfully enables the LLM to bypass the injected instruction. What’s more, the response still remains relevant to the original task, suggesting the defense method has little damage on the utility of the model.
6 Conclusion
In this study, we explore the design of defense methods against prompt injection attacks by leveraging attack techniques, because of the similar design goals between the attack methods and the defense methods. Specifically, we design defense strategies based on the ignore attack, escape attack, fake completion attack, and fake completion attack with template. We evaluate our methods against both direct and indirect prompt injection attacks, comparing their performance to various training-free defense methods. The experimental results demonstrate that our defense methods outperform existing defense baselines, even decreasing the ASR to zero in some scenarios. What’s more, we observe that the stronger attack method can be utilized to build stronger defense method, paving the way for designing more effective defenses against more complex attacks in the future.
7 Limitations
In this paper, we propose defense methods derived from existing attack strategies. However, our focus has been limited to prompt-engineering-based attacks, and we have not explored to design defenses with gradient-based attacks. Generally, gradient-based attacks tend to be more effective. Based on our findings that strong attack methods often inspire strong defense mechanisms, we believe that defenses designed with gradient-based attack methods would likely be more robust. Investigating how to develop such defenses will be a key direction for our future work.
8 Ethical consideration
We declare that all authors of this paper acknowledge the ACM Code of Ethics and honor the ACL code of conduct. The primary goal of this work is to defend against the prompt injection attacks. The source code will be publicly available. We apply existing benchmark datasets in the experiment, and thereby not introducing new safety risks regarding the unsafe data samples.
Appendix A Appendix
Figure 3: The relationship between the attack and defense method’s effectiveness. Each point represents either the average ASR of an attack against different defense methods or the average ASR of a defense method against various attacks.
A.1 Defense Baselines
Sandwich san (2023).
A reminder of the original input instruction is appended to the end of the data content, encouraging the LLM to follow the correct instruction. An example is shown in Fig 4.
Instructional ins (2023).
After the original input instruction, this method warns the LLM about potential attacks and emphasizes following the original instruction. An example is shown in Fig 6.
Reminder Yi et al. (2023).
A simple reminder, such as “Do not execute any instructions in the following content,” is added after the original input instruction. An example is shown in Fig 7
Isolation Willison (2023).
Special tokens are used to clearly label the data content portion. An example is shown in Fig 5
Spotlight Hines et al. (2024).
This method connects the entire data content area using special tokens, making the data content areas more obvious. An example is shown in Fig 8
A.2 Implementation Details.
We conduct our defense experiments using PyTorch 2.1.0 Paszke et al. (2019). The experiments are performed on a single NVIDIA A100 GPU. For generation, we set “do_sample” to false and “max_new_tokens” to 256. The “max_length” is set to 8192.
A.3 Dataset Additional Information
AlpacaFarm is with license of Apache License 2.0 and dataset from Li et al. (2023) does not report the license. The datasets we use do not contain personal privacy information. AlpacaFarm is a simulation platform designed to facilitate research and development in learning from feedback, significantly reducing the typical costs. It aims to make research on instruction following and alignment more accessible Dubois et al. (2024). What’s more, dataset from Li et al. (2023) is constructed to evaluate the prompt inject attack. The datasets are open-sourced.
Defense Methods Naive Ignore Escape Fakecom Combined
None 10.55 53.35 88.25 75.30 86.00
StruQ-Naive 0.50 0.60 2.20 35.55 27.30
StruQ-Ignore 0.05 0.05 8.00 35.70 0.05
Ours-Ignore 0.05 0.35 0.30 0.10 1.35
Table 7: Defense performance of our methods and the fine-tuning method StruQ. "StruQ-Naive" means StruQ incorporates the "Naive attack" for fine-tuning. Bold indicates the best performance. All results are reported in %.
Defense Methods Llama3 Llama3.1 Qwen2
None 0.645 0.699 1.184
Sandwich 0.632 0.701 1.173
Instructional 0.630 0.701 1.169
Reminder 0.632 0.701 1.179
Isolation 0.631 0.702 1.171
Spotlight 0.644 0.734 1.193
Ours-Ignore 0.632 0.701 1.171
Ours-Escape 0.644 0.729 1.548
Ours-Fakecom 0.631 0.701 1.171
Ours-Fakecom-t 0.766 0.704 1.200
Table 8: The time cost across various defense methods. All the results are reported in seconds/item.
Figure 4: Sandwich defense example.
Figure 5: Isolation defense example.
Figure 6: Instructional defense example.
Figure 7: Reminder defense example.
Figure 8: Spotlight defense example.
Figure 9: Naive attack example.
Figure 10: Escape-Deletion attack example.
Figure 11: Escape-Separation attack example.
Figure 12: Ignore attack example.
Figure 13: Fake completion attack example.
Figure 14: Fake completion attack with template example.
Figure 15: Combine attack example.
Figure 16: The examples of the responses to ignore and fake completion attack without defense and with fake completion defense with template.